If two blood vessels, each with a resistance of 1 (arbitrary units), were placed together in series, their total resistance would be ____; if placed in parallel, their total resistance would be _____. a. 2 ; 0.5

b. 0.5 ; 2

c. 2 ; 1

d. 1 ; 2

e. 2 ; 2
Answer #1

A. 2 AND 0.5

ANSWER:

THE FORMULA TO CALCULATE THE TOTAL RESISTANCE WHEN THEY PLACED TOGETHER:

Let as assume the total resistance as Rx, first blood vessel as R1 and second blood vessel as R2. If two blood vessels contains 1 arbitary units each, then if they placed together, the formula will be Rx= R1+R2. Therefore, Rx = 1+1 =2

THE FORMULA TO CALCULATE PARALLEL RESISTANCE OF THEIR RESISTANCE (Rx) = 1 divided by 1/R1 + 1/R2

Then it becomes, 1 divided by 1/1 + 1/1 = 1/2 = 0.5. So the parallel resistance will be 0.5.